OpenStudy (anonymous):

In how many different ways can you answer a multiple choice test that has 5 questions and 3 choices for each answer?

OpenStudy (anonymous):

3*3*3*3*3 = 3^5 = 243 ways

OpenStudy (lgbasallote):

use combinations \[nCr = \frac{n!}{r!(n-r)!}\] so in this case \[5C3 = \frac{5!}{3!(5-3)!}\] does that help?

OpenStudy (anonymous):

yes , thanks. :)

OpenStudy (lgbasallote):

welcome ^_^

OpenStudy (anonymous):

what do the exclamation marks mean?

OpenStudy (anonymous):

I dont 10 is the answer @lgbasallote

OpenStudy (anonymous):

the choices are: 15 125 243

OpenStudy (lgbasallote):

it means factorial...for example \[6! = 6\times 5 \times 4 \times 3 \times 2\times 1\] \[7! = 7 \times 6 \times 5 \times 4 \times 3 \times 2 \times 1\] \[5! = 5\times 4 \times 3 \times 2 \times 1\] and so on...

OpenStudy (anonymous):

 which is the correct answer